# Domain Specific Pruning of Large Mixture-of-Experts Models with Few-shot Demonstrations

### v1

@@ Abstract
Mixture-of-Experts (MoE) models achieve a favorable trade-off between performance and inference efficiency by activating only a subset of experts. However, the memory overhead of storing all experts remains a major limitation, especially in large-scale MoE models such as DeepSeek-R1 (671B). In this study, we investigate domain specialization and expert redundancy in large-scale MoE models and uncover a consistent behavior we term few-shot expert localization, with only a few demonstrations, the model consistently activates a sparse and stable subset of experts. Building on this observation, we propose a simple yet effective pruning framework, EASY-EP, that leverages a few domain-specific demonstrations to identify and retain only the most relevant experts. EASY-EP comprises two key components: output-aware expert importance assessment and expert-level token contribution estimation. The former evaluates the importance of each expert for the current token by considering the gating scores and magnitudes of the outputs of activated experts, while the latter assesses the contribution of tokens based on representation similarities after and before routed experts. Experiments show that our method can achieve comparable performances and 2.99×2.$9$ throughput under the same memory budget with full DeepSeek-R1 with only half the experts. Our code is available at https://github.com/RUCAIBox/EASYEP.
Figure 1: Throughput and Performance Comparison of DeepSeek-R1 on AIME2024 with Varying Expert Counts Using EASY-Prune. We deploy R1 with two 8$\times$ H800 for 224 and 256 experts while one 8×8$\times$8 × H800 is employed for others. The throughputs of latter configurations are multiplied by 2.
@@ 1 Introduction
Mixture-of-Experts (MoE) architectures have been widely adopted as the backbones of various large language models (LLMs) due to their efficiency of scaling parameters without proportional computational overhead [1, 2, 3, 4]. However, the deployment of large MoE models imposes substantial memory requirements. Taking DeepSeek-R1 (671B) [1] as an example, it takes about 1500 GB under BF16 precision and 750 GB under FP8 precision, necessitating 4×8 A800 or 2×8 H800 GPU configurations, respectively. This underscores the critical need to explore lite deployment strategies for large-scale MoE models like DeepSeek-R1.
@@
Various training-free approaches have been proposed to alleviate the inference memory demands of MoE models. Router-based pruning methods leverage statistical information from router activation and gating values to estimate expert importance [5], while perturbation-based strategies search for the optimal subsets of experts to minimize the distribution drift of hidden states [6, 7]. Despite these advancements, achieving an optimal trade-off between efficiency and performance remains a challenge. Router-based methods often fail to accurately identify critical experts, whereas perturbation-based approaches become computationally prohibitive as the number of experts scales up. Moreover, existing methods are primarily designed for MoEs with a few experts per layer (e.g., Mixtral 8$\times$7B [2][6]. This highlights the need for more scalable and accurate pruning strategies tailored to large-scale MoE models. While such a scale poses challenges for current approaches, it also brings new opportunities: the increased granularity and domain specialization of experts in models like DeepSeek-R1 [8, 9] make them especially amenable to domain-specific pruning at high compression ratios [10].
@@
In this study, we analyze how expert activations in a large MoE model vary across domains and respond to a small number of demonstration samples. We observe a consistent behavior: given just a few domain-specific examples, the model tends to activate a stable and sparse subset of experts that are highly relevant to that domain. We refer to this phenomenon as few-shot expert identification. To better understand this behavior, we first examine domain-level differences in expert activation distributions and the role of high gating-value experts. We then analyze how the number of demonstrations and the similarity across datasets within the same domain affect expert identification. Our key findings are as follows: (1) High gating-value experts are strongly domain-specific, which play a critical role in their respective domains but have minimal or no influence on unrelated tasks; (2) A small number of demonstrations is sufficient to identify these experts reliably, and the identified experts generalize well to other datasets within the same domain.
@@
Based on our observations, we introduce a domain-specific pruning framework that leverages few-shot demonstrations to address memory constraints in large-scale MoE models. Our approach begins by sampling a small number of task demonstrations from a specific domain and generating responses with the original model, serving as a calibration set. We then introduce a pruning method, Expert Assessment with Simple Yet-effective scoring for Expert Pruning, a.k.a., EASY-EP. It aims to quantify expert importance and retain a fixed-size subset of top-scoring experts. This proposed metric consists of two complementary components: output-aware expert importance assessment and expert-level token contribution estimation. The former component jointly considers gating values and L2 Norm of expert outputs to evaluate per-token expert importance, while the latter component uses similarity between input and residual-connected output of routed experts to measure the contribution of individual tokens to the overall expert score. Thus, a single forward pass is sufficient to identify target domain-specific experts without the need for backpropagation or repeated forward evaluations.
@@
To evaluate the efficacy of our approach, we conducted systematic experiments on DeepSeek-R1 using six benchmark datasets spanning math, coding, science, and agent execution capabilities. Specifically, our method achieves $100.38\%$ and $95.10\%$ performance compared to the full model while retaining only $50\%$ of experts on domain-specific and mixed-domain pruning settings, respectively. As the compression ratio increases, the performances degrade slowly, indicating robustness to pruning. Additionally, under identical memory constraints, pruning $50\%$ of the experts leads to a 2.99$\times$ increase in inference throughput under the setting of 1K input and 1K output lengths, highlighting the practical utility of our framework for real-world deployments.
@@ 2 Background
Recently, Mixture-of-Experts (MoE) architectures have been widely adopted as the backbone of LLMs due to their ability to scale parameters efficiently without incurring additional computational costs [2, 10]. This architecture introduces MoE modules where parameters are dynamically activated to replace traditional feedforward networks (FFNs) In the $l$-th layer, a MoE module consists of a router $\operatorname{G}(\cdot)$, $N$ routed experts $\{\operatorname{E}^{l}_{1}(\cdot),\dots,\operatorname{E}^{l}_{N}(\cdot)\}$, and an optional shared expert $\operatorname{E}^{l}_{s}(\cdot)$  [^1] Shared experts were introduced by the DeepSeekMoE series [10, 1, 11, 12], but are not considered in our analysis and method. at layer $l$. Given an input representation sequence $\mathbf{H}^{l}=\{\bm{h}^{l}_{1},\dots,\bm{h}^{l}_{T}\},\forall\bm{h}_{t}^{l}\in\mathbb{R}^{D}$, the router computes the logits of each routed expert for the $t$-th token and applies a gating function on the Top-$K$ logits to obtain the gating values:
$g^{l}_{i,t}=\begin{cases}\operatorname{G}^{l}_{i}(\bm{h}_{t}^{l}),\quad&\operatorname{G}^{l}_{i}(\bm{h}_{t}^{l})\in\text{Top-}K({\operatorname{G}^{l}_{1}(\bm{h}_{t}^{l}),\dots,\operatorname{G}^{l}_{N}(\bm{h}_{t}^{l})}),\\ 0,&\text{otherwise}.\end{cases}$ (1)
Subsequently, the Top-$K$ selected experts are activated, and their outputs are aggregated via weighted summation. The final output is obtained by summing the input hidden states and the output of routed experts as follows:
$\tilde{\bm{h}}^{l}_{t}={\bm{h}}^{l}_{t}+\sum_{i=1}^{N}g^{l}_{i,t}\cdot\operatorname{E}_{i}^{l}(\bm{h}^{l}_{t})$ (2)
With the gating values of the router, we define two metrics to assess the importance of each expert, i.e., frequency and gating scores [5]. For all tokens in a calibration set and an expert $\operatorname{E}_{i}^{l}$, we define frequency $f^{l}_{i}$ as the number of times each expert is activated, while gating scores $r^{l}_{i}$ is defined as the total sum of the gating values when each expert is activated.
@@
where $M$ is the number of demonstrations in the calibration set, and $T_{n}$ denotes the number of tokens per demonstration.
@@ 3 Empirical Analysis of Experts
Previous work has demonstrated that the expert distributions of MoEs with few experts (e.g., Mixtral $8\times 7$B) mainly depend on the syntax structures instead of tasks and domains of input data [2]. However, the large-size MoE models are equipped with various fine-grained experts, e.g., DeepSeek-R1 has 256 experts in each layer. Thus, the experts may be more specialized and store distinct knowledge and capacities in their parameters. To investigate this, we empirically investigate what we refer to as few-shot expert identification, a phenomenon in which domain-specific experts can be reliably identified using only a small number of demonstrations.
Figure 2: Overlap ratios of experts with top gating scores on different datasets.
@@ 3.1 Expert Specialization Across Domains
To investigate whether experts in large MoE models exhibit domain-specific specialization, we select AIME-2023  [^2] https://huggingface.co/datasets/di-zhang-fdu/AIME_1983_2024, GPQA-main [13], and LiveCodeBench-V3 [14] as representative datasets for the domains of math, science, and coding, respectively. We conduct experiments using the representative model DeepSeek-R1 on these datasets and extract the gating scores across different domains (as described in Section 2). By analyzing the expert activation distributions across these domains, we aim to reveal how expert utilization varies and assess the degree of specialization.
Distinct Expert Distribution Across Domains. We first rank all experts at each layer by their gating scores and select the top-$M$ experts for each dataset, where $M\in\{2,16,128\}$. We then measure the overlap of top-ranked experts across different domains and visualize how this overlap varies with $M$ in Figure 2. We observe that the top-2 experts are completely disjoint across datasets, and even as $M$ increases, a substantial number of non-overlapping experts remain. This indicates that large MoE models contain domain-specialized experts that are predominantly activated in their respective domains.
Table 1: Results of removing domain-specific experts.
@@
Full 78.00 70.91 63.32
Math 67.33 (-10.67) 69.19 (-1.72) 65.27 (+1.95)
Code 78.67 (+0.67) 71.72 (+0.81) 55.68 (-6.64)
Science 79.33 (+1.33) 59.09 (-11.82) 61.07 (-2.25)
Impact of Removing Domain-Specific Experts. In order to explore the importance of domain-specific experts, we remove those that appear among the top-128 by gating score in one domain but not in the top-128 of any other domain. We then evaluate each pruned model on tasks from the same domain as the calibration data (i.e., in-domain), as well as on tasks from other domains (i.e., out-of-domain), to assess the generalizability of the retained experts. As shown in Table 1, pruning these experts leads to significant performance degradation on in-domain tasks while having minimal impact on out-of-domain tasks. From these findings, we can conclude that domain-specific experts play a critical role in relevant domain but are redundant for other domains.
@@ 3.2 Expert Locality Within One Domain
Beyond examining the expert specialization across different domains, we further investigate the distribution of experts within the same domain. Specifically, we study how the number of demonstrations affects pruning and whether the identified critical experts remain highly activated on other datasets within the same domain.
Figure 3: Left: Overlap ratio of top-128 experts with different number of demonstrations. Right: Overlap ratio of top-128 experts pruned with different math datasets.
Effect of Calibration Set Size. Given the presence of domain-specific experts in DeepSeek-R1, a fundamental question arises: How many demonstrations are necessary to accurately identify these key experts? To investigate this, we sample varying numbers of demonstrations (i.e., 1,3,5,10,25,100) from LiveCodeBench-v3 [14]. For each setting, we compute the average gating scores and retain the top 128 experts. Based on these selections, we then calculate the pairwise overlap ratios between expert sets derived from different demonstration sizes. As illustrated in Figure 3 (Left), even with just five demonstrations, over $90\%$ of the critical experts can be effectively identified. Furthermore, 25 demonstrations are sufficient to capture all domain-specific experts ($99\%$), with additional demonstrations yielding only marginal improvements. These results underscore the feasibility of few-shot domain-specific expert pruning.
Consistency of Expert Activation Across Datasets. To investigate the consistency of domain-specific expert activation across datasets within the same domain, we conducted experiments with DeepSeek-R1 on four math datasets: AIME-2023, AIME-2024, AIME-2025, and HMMT-Feb 2025. Specifically, we perform expert pruning under the 25-shot setting for each dataset and retain the top 128 experts based on their average gating scores. We then compute the pairwise overlap ratios between each pair of datasets. As shown in Figure 3 (Right), the overlaps exceed 84% across all math datasets, revealing a high degree of consistency in domain-specific expert activation. This indicates that domain-specific expert activation patterns are largely transferable within the same domain. Despite some dataset-specific differences, the overall expert overlap remains strong and stable.
@@ 4.1 Overview
Motivated by our earlier observation of few-shot expert identification phenomena in Section 3, we propose an expert pruning framework to reduce memory costs. Given a small set of demonstrations from the target domain, we run them through the MoE model and collect statistics on expert activation, considering both the inputs and outputs of each expert for pruning purposes. To effectively identify domain-specific experts in large MoE models, we introduce a simple yet effective expert pruning method, EASY-EP. Specifically, we first compute the output-aware expert importance $c_{i,t}^{l}$ as the product of the expert output L2 norm and its corresponding gating value.
Next, we determine the expert-level token contribution $s^{l}_{t}$ based on representation similarity between before and after expert computation. The final expert score $I(\operatorname{E}_{i}^{l})$ is obtained by multiplying these two terms:
@@ 4.2 Output-Aware Expert Importance Assessment
To assess the importance of each expert, prior router-based expert pruning methods [5] assume activated expert gating scores can reflect their importance. However, this assumption has not been thoroughly examined. To analyze this, we begin by examining the aggregated output $\bar{\bm{h}}_{t}^{l}$ from all routed experts for a given token:
@@
This indicates that the each expert’s contribution to the final output is bounded by the product of its gating value and the L2 norm of its output, $g_{i,t}^{l}\lVert\bm{e}_{i,t}^{l}\rVert$. To empirically validate this insight, we visualize in Figure 5 (Left) the relationship between gating scores $g_{i,t}^{l}$ and the product $g_{i,t}^{l}\lVert\bm{e}_{i,t}^{l}\rVert$. We can observe an expert with a large gating score may still produce outputs with low L2 Norm, ultimately resulting in a limited influence on the final output. Therefore, instead of using only the gating value, we define the importance of an expert for a given token as the product of its gating value and L2 Norm of output, as formalized in the following equation:

### v2

@@ Abstract
Mixture-of-Experts (MoE) models achieve a favorable trade-off between performance and inference efficiency by activating only a subset of experts. However, the memory overhead of storing all experts remains a major limitation, especially in large-scale MoE models such as DeepSeek-R1 (671B). In this study, we investigate domain specialization and expert redundancy in large-scale MoE models and uncover a consistent behavior we term few-shot expert localization, with only a few in-domain demonstrations, the model consistently activates a sparse and stable subset of experts on tasks within the same domain. Building on this observation, we propose a simple yet effective pruning framework, EASY-EP, that leverages a few domain-specific demonstrations to identify and retain only the most relevant experts. EASY-EP comprises two key components: output-aware expert importance assessment and expert-level token contribution estimation. The former evaluates the importance of each expert for the current token by considering the gating scores and L2 norm of the outputs of activated experts, while the latter assesses the contribution of tokens based on representation similarities before and after routed experts. Experiments on DeepSeek-R1 and DeepSeek-V3-0324 show that our method can achieve comparable performances and 2.99×2.$9$ throughput under the same memory budget with full model with only half the experts.
@@ 1 Introduction
Mixture-of-Experts (MoE) architectures have been widely adopted as the backbones of various large language models (LLMs) due to their efficiency of scaling parameters without proportional computational overhead (1; 2; 3; 4). However, the deployment of large MoE models imposes substantial memory requirements. Taking DeepSeek-R1 (671B) (1) as an example, it takes about 1500 GB under BF16 precision and 750 GB under FP8 precision, necessitating 4×8 A800 or 2×8 H800 GPU configurations, respectively. This underscores the critical need to explore lite deployment strategies for large-scale MoE models like DeepSeek-R1.
Figure 1: Throughput and performance comparison of DeepSeek-R1 on AIME2024 with varying expert numbers using EASY-EP. We deploy DeepSeek-R1 with two 8$\times$ H800 for 224 and 256 experts, while one 8×8$\times$8 × H800 for others. The throughputs of the latter configurations are multiplied by 2.
@@
Various training-free approaches have been proposed to alleviate the inference memory demands of MoE models. Expert pruning reduces memory by removing less important experts. Among them, router-based methods use expert activation statistics to estimate importance (5), while perturbation-based ones select expert subsets to minimize hidden state drift (6; 7). However, the former may fail to identify key experts, and the latter is expensive as the number of experts increases. Expert merging combines similar experts to reduce their number (8; 9), but can cause neuron misalignment of experts in MoE models trained from scratch (10; 11). Moreover, existing methods are primarily designed for MoEs with a few experts per layer (e.g., Mixtral 8$\times$7B (2)(6). This highlights the need for more scalable and accurate pruning strategies tailored to large-scale MoE models. While such a scale poses challenges for current approaches, it also brings new opportunities: the increased granularity and domain specialization of experts in models like DeepSeek-R1 (12; 13) make them especially amenable to domain-specific pruning at high compression ratios (14).
@@
In this study, we analyze how expert activations in a large MoE model vary across domains and respond to a small number of demonstration samples. We observe a consistent behavior: given just a few domain-specific examples, the model tends to activate a stable and sparse subset of experts that are highly relevant to the target domain. We refer to this phenomenon as few-shot expert localization. To better understand the mechanism behind this behavior, we focus on two factors: the domain specialization of experts and the sufficiency of limited demonstrations. Our key findings are as follows: (1) High gating-value experts are strongly domain-specific, consistently dominating activations within their respective domain while remaining inactive in unrelated ones. (2) A small number of demonstrations suffices to reliably trigger these experts, and they generalize well to other unseen datasets within the same domain.
@@
Based on our observations, we introduce a domain-specific pruning framework that leverages few-shot demonstrations to address memory constraints in large-scale MoE models. Our approach begins by sampling a small number of task demonstrations from a specific domain and generating responses with the original model, serving as a calibration set. To identify and retain the most critical experts, we then develop a pruning method, Expert Assessment with Simple Yet-effective scoring for Expert Pruning, a.k.a., EASY-EP. This method estimates expert importance and retain a fixed-size subset of top-scoring experts. EASY-EP consists of two complementary components: (1) output-aware expert importance assessment, which combines gating values and the L2 norm of expert outputs to estimate per-token expert importance, and (2) expert-level token contribution estimation, which measures the similarity between the input and the residual-connected output of routed experts to measure each token’s contribution to the overall expert score. Notably, the entire scoring process requires only a single forward pass, eliminating the need for backpropagation or repeated evaluations.
@@
To evaluate the efficacy of our approach, we conducted systematic experiments on DeepSeek-R1 and DeepSeek-V3-0324 using eight benchmark datasets covering math, coding, science, finance, medicine, and agent execution capabilities. Specifically, with only retraining $50\%$ experts, our method can keep comparable performances under the domain-specific pruning settings while achieving over $90\%$ of the full model’s performances on DeepSeek-R1 and even better performances on DeepSeek-V3-0324 under the mixed-domain pruning settings. As shown in Figure 1, the performances degrade only slowly with the increase of compression ratio, indicating robustness to pruning. Additionally, under identical memory constraints, pruning $50\%$ of the experts yields a 2.99$\times$ increase in inference throughput for sequences of 1K input and 1K output lengths, highlighting the practical utility of our framework in real-world deployments.
@@ 2 Background
MoE architectures introduce MoE modules where parameters are dynamically activated to replace feedforward networks (FFNs) (2; 1). Specially, in the $l$-th layer, a MoE module contains a router $\operatorname{G}(\cdot)$ and $N$ routed experts $\{\operatorname{E}^{l}_{1}(\cdot),\dots,\operatorname{E}^{l}_{N}(\cdot)\}$  [^1] Shared experts are employed in some MoE models (1; 14; 15; 16), but are not considered in this work.. Given an input representation sequence $\mathbf{H}^{l}=\{\bm{h}^{l}_{1},\dots,\bm{h}^{l}_{T}\},\forall\bm{h}_{t}^{l}\in\mathbb{R}^{D}$, the router computes the logit of each expert for the $t$-th token and applies a gating function on the Top-$K$ logits to obtain the gating values $g_{i,t}^{l}$ (the gating values of deactivated experts $g_{i,t}=0$). The Top-$K$ experts are activated, and their outputs are aggregated via weighted summation. The final output is obtained by residual connections of the input and output of experts:
$\bar{\bm{h}}^{l}_{t}=\sum_{i=1}^{N}g^{l}_{i,t}\cdot\operatorname{E}_{i}^{l}(\bm{h}^{l}_{t}),~{}~{}~{}~{}\tilde{\bm{h}}^{l}_{t}={\bm{h}}^{l}_{t}+\bar{\bm{h}}_{t}^{l}$ (1)
With the gating values of the router, we define two metrics to assess the importance of each expert, i.e., frequency and gating scores (5). For all tokens in a calibration set and an expert $\operatorname{E}_{i}^{l}$, we define the frequency $f^{l}_{i}$ as the number of times each expert is activated, while the gating scores $r^{l}_{i}$ is defined as the total sum of the gating values when each expert is activated, as shown in Equation 2. Here, $M$ is the size of the calibration set, and $T_{n}$ denotes the number of tokens per demonstration.
@@ 3 Empirical Analysis of Experts
Previous work has demonstrated that the expert distributions of MoEs with few experts (e.g., Mixtral $8\times 7$B) mainly depend on the syntax structures instead of domains (2). However, recent large-size MoE models are equipped with various fine-grained experts (e.g., DeepSeek-R1 has 256 experts per layer), which may be more specialized and store distinct knowledge and capacities in their parameters. Motivated by this, we empirically study a phenomenon we term few-shot expert localization, where domain-specific experts can be reliably identified using only a handful of demonstrations.
Figure 2: (A), (B): Overlap ratios of experts with top gating scores on different datasets. (C): Overlap ratio of top-128 experts with different numbers of demonstrations. (D): Overlap ratio of top-128 experts pruned with different math datasets.
@@ 3.1 Expert Specialization Across Domains
To assess whether experts in large MoE models exhibit domain-specific specialization, we select AIME-2023 (17), GPQA-main (18), and LiveCodeBench-V3 (19) as calibration datasets for the domains of math, science, and coding, respectively. We conduct experiments using the representative model DeepSeek-R1 on these datasets and extract the gating scores across different domains (as described in Section 2). By analyzing the expert activation distributions across these domains, we aim to reveal how expert utilization varies and assess the degree of specialization.
Distinct Expert Distribution Across Domains. We first rank all experts at each layer by their gating scores and select the top-$16$ and top-$128$ experts for each dataset. Subsequently, we measure the overlap of top-ranked experts across different domains and visualize the overlap ratios in Figure 2, where (A) corresponds to top-16 and (B) to top-128. We observe that the top-16 experts are largely disjoint across datasets. When expanding the selection to top-$128$, the degree of overlap increases, but a significant portion of the experts remains domain-specific. This indicates that large MoE models contain domain-specialized experts that are predominantly activated in their respective domains. We show experiments on different numbers of top experts and layer variations in Appendix B.
Table 1: Results of removing domain-specific experts. Bold denotes in-domain results.
@@
Full 77.08 70.91 63.32
Math 67.33 (-9.75) 69.19 (-1.72) 65.27 (+1.95)
Code 78.67 (+1.59) 71.72 (+0.81) 55.68 (-6.64)
Science 79.33 (+2.25) 59.09 (-11.82) 61.07 (-2.25)
Impact of Removing Domain-Specific Experts. In order to explore the importance of domain-specific experts, we remove those that appear in the top-128 (by gating score) in a specific domain but not in any others. We then evaluate each pruned model on tasks from the same domain as the calibration data (i.e., in-domain), and on tasks from other domains (i.e., out-of-domain), to assess the generalization behavior of the remaining experts. As shown in Table 1, pruning these experts leads to significant performance degradation on in-domain tasks while having minimal impact on out-of-domain tasks. These results suggest that domain-specific experts play a critical role in the relevant domain but are redundant for other domains.
@@ 3.2 Expert Locality Within One Domain
Beyond examining the expert specialization across different domains, we examine the locality and stability of expert activation within a single domain. We investigate how the number of demonstrations and the choice of calibration set influence expert selection patterns under the same domain setting.
Effect of Calibration Set Size. Given the presence of domain-specific experts in DeepSeek-R1, a fundamental question arises: How many demonstrations are necessary to accurately identify these key experts? To answer this, we sample varying numbers of demonstrations (i.e., 1, 5, 25, 100) from LiveCodeBench-v3 (19). For each setting, we compute the average gating scores and retain the top 128 experts. Based on these selections, we then calculate the pairwise overlap ratios between expert sets derived from different demonstration sizes. As illustrated in Figure 2 (C), even with just five demonstrations, over $90\%$ of the critical experts can be effectively identified. Furthermore, 25 demonstrations are sufficient to capture all domain-specific experts ($99\%$), with additional demonstrations yielding only marginal improvements. These results underscore the feasibility of few-shot domain-specific expert pruning.
Consistency of Expert Activation Across Datasets. To investigate the consistency of domain-specific experts across datasets within the same domain, we conduct experiments with DeepSeek-R1 on four math datasets: AIME-2023, AIME-2024, AIME-2025, and HMMT-Feb 2025 (17). Specifically, we perform expert pruning under the 25-shot setting for each dataset and retain Top-128 experts based on their average gating scores. We then compute the pairwise overlap ratios between each pair of datasets. As shown in Figure 2 (D), the overlaps exceed $84\%$ across all math datasets, revealing a high degree of consistency in domain-specific expert activation. This indicates that domain-specific expert activation patterns are largely transferable within the same domain. Despite some dataset-specific differences, the overall expert overlap remains strong and stable.
Empirical analysis with another metric is shown in Appendix C.
@@ 4.1 Overview
Motivated by our earlier observation of few-shot expert localization phenomena in Section 3, we propose an expert pruning framework to reduce memory costs. We use a small set of target-domain demonstrations to run the MoE model and collect expert activation statistics, considering both inputs and outputs for pruning. To effectively identify domain-specific experts in large MoE models, we introduce a simple yet effective expert pruning method, EASY-EP. Specifically, we first compute the product of the expert output L2 norm and its corresponding gating value as the output-aware expert importance $c_{i,t}^{l}$. Next, we determine the expert-level token contribution $s^{l}_{t}$ based on the similarity of representations before and after expert computation. The final expert score $I(\operatorname{E}_{i}^{l})$ is obtained by aggregating the product of two terms over all tokens:
@@ 4.2 Output-Aware Expert Importance Assessment
To assess the importance of each expert, prior router-based expert pruning methods assume activated expert gating scores can reflect their importance (5). However, this assumption has not considered the influence of experts. To further assess the contribution of each routed expert, we example the aggregated output $\bar{\bm{h}}_{t}^{l}$ from all routed experts for a given token:
@@
This indicates that each expert’s contribution to the final output is bounded by the product of its gating value and the L2 norm of its output, $g_{i,t}^{l}\lVert\bm{e}_{i,t}^{l}\rVert$. An expert with a large gating score may still produce outputs with low L2 norm, ultimately resulting in a limited influence on the final output, which are empirically verified in Appendix D.1. Therefore, instead of using only the gating value, we define the importance of an expert for a given token as the product of its gating value and L2 norm of output, as formalized in the following equation: